# Insights into the Lottery Ticket Hypothesis and Iterative Magnitude Pruning

### v2

@@ I Introduction
Many research works related to the lottery ticket hypothesis have been published since its advent to demystify the mechanisms and principles governing it. Frankle et al. [7] studied the relationship between the percentage of weights remaining and interpretability [^1] Ability to understand and explain how the network arrived at its decisions. of a network in case of magnitude pruning. They demonstrated that pruning does not harm interpretability until very few parameters remain in the network and there is an accuracy drop. The authors demonstrated that the network parameters considered superfluous for accuracy by the pruning algorithm are also superfluous for interpretability. However, pruning does not make the network more interpretable either. The network dissection technique given by Bau et al. [8] was used to measure the interpretability of the pruned network. In another work, Frankle et al. [9] demonstrated that the pruning methods that prune at initialization (prior to training) perform better than random pruning, but their accuracy remains smaller than the magnitude pruning after training. They also demonstrated that randomly shuffling the weights pruned by these techniques within each layer preserves or improves the accuracy. Frankle et al. [10] performed instability analysis towards Stochastic Gradient Descent (SGD) noise to explain the success and failure of IMP. They demonstrated that IMP finds matching sub-networks only when they are stable to SGD noise (not affected by different data orders), and this happens at initialization in simple tasks but only after a few training iterations in complicated tasks. That is why pre-training the dense network for a few iterations (to obtain the rewind point) and rewinding the unpruned weights of the pruned network to their values at the rewind point is required in IMP. However, rewinding alone does not account for the success of IMP since randomly pruning the network and then rewinding to the rewind point does not produce matching sub-networks [10]. Hence, there is some information encoded in the associated pruning procedure, in addition to the information conveyed by the specific initialization proposed by the lottery ticket hypothesis, that leads to the success of IMP. Larsen et al. [11] demonstrated that the number of parameters required for training a deep neural network decreases as the initial loss decreases. This explains why rewinding to a later point in training leads to better accuracy. Zhang et al. [12] demonstrated the structural importance of winning tickets by showing that pruning enlarges the convex region near a desirable model. They also showed that the number of training samples required for achieving zero generalization error is proportional to the number of unpruned weights.
@@
Contributions. We perform extensive experimentation on two widely used networks, ResNet-20 and VGG-16, on a benchmark dataset, CIFAR-10. The contributions are as follows:
@@
    We demonstrate that there exist special type of minima in the loss landscape, which generalize well but have a very small volume in the original space, and the IMP procedure exposes such minima, which otherwise remain hidden.
@@ III Methodology
In this section, we limit ourselves to a top-level description of the experiments conducted for the study. Some of the finer details are relegated to Appendix I.
We perform experimentation on two widely used networks, ResNet-20 and VGG-16, on a benchmark dataset, CIFAR-10. We use iterative magnitude pruning with weight rewinding (IMP-WR) and run $10$ iterations of IMP in the case of ResNet-20 and $12$ iterations of IMP in the case of VGG-16. We prune 20% of the smallest magnitude weights in each iteration. The unpruned weights are rewound to their values at $2000^{th}$ training step of the original dense network in each iteration. The main part of the manuscript contains the results obtained on ResNet-20. The corresponding results for VGG-16 are given in Appendix II. For an initial and quick appreciation, a plot of training loss and test accuracy at different levels of IMP-WR in the case of ResNet-20 is presented in Fig. 1.
@@
One-shot pruned network is obtained by pruning the weights of the trained dense network $W_{(0)}^{(min\_(0))}$ based on magnitude pruning in one go to attain the desired sparsity [^5] Same sparsity as that of $W_{(10)}^{(min\_(10))}$ in case of ResNet-20, and same sparsity as that of $W_{(12)}^{(min\_(12))}$ in the case of VGG-16. and then rewinding the unpruned weights to their values at $W^{(rewind\_point)}$ and retraining. We represent the solution obtained using one-shot pruning by $W^{(one\_shot)}$.
Fine-tuned network is obtained by pruning $20\%$ smallest magnitude weights from $W_{(9)}^{(min\_(9))}$ in case of ResNet-20 and by pruning $40\%$ smallest magnitude weights from $W_{(11)}^{(min\_(11))}$ in the case of VGG-16 and then re-training the unpruned weights (without rewinding) with a learning rate of $0.001$ for $40$ epochs. We represent the solution obtained using fine-tuning by $W^{(FT)}$.
Randomly initialized pruned network is obtained by pruning $20\%$ smallest magnitude weights from $W_{(9)}^{(min\_(9))}$ in case of ResNet-20 and by pruning $40\%$ smallest magnitude weights from $W_{(11)}^{(min\_(11))}$ in the case of VGG-16 and then randomly initializing the unpruned weights and retraining. We call this network Randomly Initialized Pruned Network and represent the solution by $W^{(RIPN)}$.
Randomly pruned network is obtained by randomly pruning $20\%$ weights from $W_{(9)}^{(min\_(9))}$ and then rewinding and retraining in case of ResNet-20. We call this network Randomly Pruned Network and represent the solution by $W^{(RPN\_1)}$. We also consider another randomly pruned network in the case of ResNet-20 by randomly pruning the weights of $W_{(0)}^{(min\_(0))}$ in one go to attain the desired sparsity and then rewinding and re-training. We represent the solution obtained in this manner by $W^{(RPN\_2)}$. Note that the only difference between $W^{(RPN\_2)}$ and $W^{(one\_shot)}$ is that while obtaining $W^{(RPN\_2)}$, the weights of $W_{(0)}^{(min\_(0))}$ are pruned randomly whereas while obtaining $W^{(one\_shot)}$, the weights of $W_{(0)}^{(min\_(0))}$ are pruned using magnitude based pruning.
In the case of VGG-16, a randomly pruned network has been obtained by randomly pruning $40\%$ weights from $W_{(11)}^{(min\_(11))}$ and then rewinding and retraining. We represent the solution obtained in the aforementioned manner by $W^{(RPN)}$.
A comparison of training loss and test accuracy between $W_{(10)}^{(min\_(10))}$, $W^{(one\_shot)}$, $W^{(FT)}$, $W^{(RIPN)}$, $W^{(RPN\_1)}$ and $W^{(RPN\_2)}$ in case of ResNet-20 presented in Fig. 2 shows that $W_{(10)}^{(min\_(10))}$ outperforms the other solutions. To explain the reason behind this, we study the loss landscape of these networks.
Figure 2: Comparison of training loss and test accuracy between $W_{(10)}^{(min\_(10))}$, $W^{(one\_shot)}$, $W^{(FT)}$, $W^{(RIPN)}$, $W^{(RPN\_1)}$ and $W^{(RPN\_2)}$. Left: Training loss. Right: Test accuracy.
@@ IV Results and Findings
IV-A Result 1: Good Minima with Small volume exist.
@@
We demonstrate that there also exist another kind of minima which have good generalization performance but have a (relatively) small volume. This implies that volume is not the only criterion for generalization performance; there is more to it. A careful experimental study leads us to hypothesize that the small volume of these minima is due to very sharp curvature in certain dimensions, but the coefficients in these dimensions are zero. Their volume measure tends to increase when these inferior dimensions are removed (possibly via pruning at another point) but is small when considered in the original space. This makes these minima almost undiscoverable by SGD in the original space but can be readily discovered in the pruned space. This is the main result of our study.
@@
Despite this at level $(L)$, SGD converges to $W_{(L)}^{(min\_(L))}$ and not to $W^{Pr{(min\_(L-1))}}_{(L)}$.
In order to find the underlying reason, we plot the trajectory of SGD for level $L$ and that for level $(L-1)$ projected on level $L$ in Fig. 4.
Figure 4: Comparison of logarithm of training loss versus epoch between level $(L)$ and level $(L-1)$ projected on level $(L)$.
@@
TABLE I: Comparison of Maximum eigen value, $\lambda_{max}$ and inverse volume of basin, $V^{{}^{\prime}}(100)$ at $W^{Pr{(min\_(L-1))}}_{(L)}$ and $W_{(L)}^{(min\_(L))}$.
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 1416.091 409.242
$W_{(1)}^{(min\_(1))}$ 1590.829 407.417
2 $W^{Pr{(min\_(1))}}_{(2)}$ 1512.426 402.678
$W_{(2)}^{(min\_(2))}$ 1349.951 399.680
3 $W^{Pr{(min\_(2))}}_{(3)}$ 1401.285 400.855
$W_{(3)}^{(min\_(3))}$ 1194.164 386.948
4 $W^{Pr{(min\_(3))}}_{(4)}$ 1266.607 392.829
$W_{(4)}^{(min\_(4))}$ 1129.884 375.590
5 $W^{Pr{(min\_(4))}}_{(5)}$ 17779.826 417.984
$W_{(5)}^{(min\_(5))}$ 1094.078 372.753
6 $W^{Pr{(min\_(5))}}_{(6)}$ 1590.785 420.431
$W_{(6)}^{(min\_(6))}$ 1065.735 373.472
7 $W^{Pr{(min\_(6))}}_{(7)}$ 2368.243 446.758
$W_{(7)}^{(min\_(7))}$ 1065.881 375.401
8 $W^{Pr{(min\_(7))}}_{(8)}$ 1882.278 432.505
$W_{(8)}^{(min\_(8))}$ 1200.277 382.051
9 $W^{Pr{(min\_(8))}}_{(9)}$ 3753.991 488.004
$W_{(9)}^{(min\_(9))}$ 1334.666 394.166
10 $W^{Pr{(min\_(9))}}_{(10)}$ 5426.848 525.621
$W_{(10)}^{(min\_(10))}$ 1575.910 408.810
[…]
@@
TABLE III: Comparison of $\lambda_{max}$ and $V^{{}^{\prime}}(100)$ at $W^{RPr{(min\_(L))}}_{(L-1)}$ and $W_{(L-1)}^{(min\_(L-1))}$.
$L$ Solution $\lambda_{max}$ $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 1742.268 418.034
$W_{(0)}^{(min\_(0))}$ 1575.277 415.42
2 $W^{RPr{(min\_(2))}}_{(1)}$ 1572.214 415.826
$W_{(1)}^{(min\_(1))}$ 1590.829 407.417
3 $W^{RPr{(min\_(3))}}_{(2)}$ 1496.211 403.674
$W_{(2)}^{(min\_(2))}$ 1349.951 399.68
4 $W^{RPr{(min\_(4))}}_{(3)}$ 1326.676 394.347
$W_{(3)}^{(min\_(3))}$ 1194.164 386.948
5 $W^{RPr{(min\_(5))}}_{(4)}$ 1291.157 391.093
$W_{(4)}^{(min\_(4))}$ 1129.884 375.59
6 $W^{RPr{(min\_(6))}}_{(5)}$ 1316.275 391.471
$W_{(5)}^{(min\_(5))}$ 1094.078 372.753
7 $W^{RPr{(min\_(7))}}_{(6)}$ 1418.483 394.086
$W_{(6)}^{(min\_(6))}$ 1065.735 373.472
8 $W^{RPr{(min\_(8))}}_{(7)}$ 1464.987 398.559
$W_{(7)}^{(min\_(7))}$ 1065.881 375.401
9 $W^{RPr{(min\_(9))}}_{(8)}$ 1619.784 411.997
$W_{(8)}^{(min\_(8))}$ 1200.277 382.051
10 $W^{RPr{(min\_(10))}}_{(9)}$ 1913.214 427.475
$W_{(9)}^{(min\_(9))}$ 1334.66 394.166
This is confirmed by a comparison of radii of the basin around $W_{(9)}^{(min\_(9))}$ and $W^{RPr{(min\_(10))}}_{(9)}$ using the method given by Huang et al. [26]. Here, the normalized cut-off value for the radii of the basins turns out to be $0.1$. A comparison of radii of the basin around $W_{(9)}^{(min\_(9))}$ and $W^{RPr{(min\_(10))}}_{(9)}$ is given in Fig. 8, and the average radii of the two basins are given in Table IV. It is clear from the figure and the table that the average radius of the basin for $W_{(9)}^{(min\_(9))}$ is larger than that for $W^{RPr{(min\_(10))}}_{(9)}$.
To sum up the discussion above, SGD converges to $W_{(L)}^{(min\_(L))}$ at level $(L)$ instead of $W^{Pr{(min\_(L-1))}}_{(L)}$ because the volume of the basin around $W_{(L)}^{(min\_(L))}$ is much larger than the volume of the basin around $W^{Pr{(min\_(L-1))}}_{(L)}$, and the path to $W_{(L)}^{(min\_(L))}$ is steeper than the path to $W^{Pr{(min\_(L-1))}}_{(L)}$. However, in the $(L-1)$ space, the volume of the basin around $W^{RPr{(min\_(L))}}_{(L-1)}$ is much smaller than the volume of basin around $W_{(L-1)}^{(min\_(L-1))}$.

### v3

@@ I Introduction
Many research works related to the lottery ticket hypothesis have been published since its advent to demystify the mechanisms and principles governing it. Frankle et al. [7] studied the relationship between the percentage of weights remaining and interpretability [^1] Ability to understand and explain how the network arrived at its decisions. of a network in case of magnitude pruning. They demonstrated that pruning does not harm interpretability until very few parameters remain in the network and there is an accuracy drop. The authors demonstrated that the network parameters considered superfluous for accuracy by the pruning algorithm are also superfluous for interpretability. However, pruning does not make the network more interpretable either. The network dissection technique given by Bau et al. [8] was used to measure the interpretability of the pruned network. In another work, Frankle et al. [9] demonstrated that the pruning methods that prune at initialization (prior to training) perform better than random pruning but their accuracy remains smaller than the magnitude pruning after training. They also demonstrated that randomly shuffling the weights pruned by these techniques within each layer preserves or improves the accuracy. Frankle et al. [10] performed instability analysis towards Stochastic Gradient Descent (SGD) noise to explain the success and failure of IMP. They demonstrated that IMP finds matching sub-networks only when they are stable to SGD noise (not affected by different data orders), and this happens at initialization in simple tasks but only after a few training iterations in complicated tasks. That is why pre-training the dense network for a few iterations (to obtain the rewind point) and rewinding the unpruned weights of the pruned network to their values at the rewind point is required in IMP. However, rewinding alone does not account for the success of IMP since randomly pruning the network and then rewinding to the rewind point does not produce matching sub-networks [10]. Hence, there is some information encoded in the associated pruning procedure, in addition to the information conveyed by the specific initialization proposed by the lottery ticket hypothesis, that leads to the success of IMP. Larsen et al. [11] demonstrated that the number of parameters required for training a deep neural network decreases as the initial loss decreases. This explains why rewinding to a later point in training leads to better accuracy. Zhang et al. [12] presented a theoretical analysis of the geometrical structure of the loss function in the case of pruned networks. They demonstrated that pruning enlarges the convex region near the optimal solution, and this geometric advantage makes pruned networks generalize better.
@@
Contributions. We perform extensive experimentation on a widely used network, ResNet-20, on a benchmark dataset, CIFAR-10. The contributions are as follows:
@@
    We demonstrate that there exist special type of solutions in the loss landscape, which generalize well but have a very small volume in the original space, and the IMP procedure exposes such solutions, which otherwise remain hidden.
@@ III Methodology
We perform experimentation on a widely used network, ResNet-20, on a benchmark dataset, CIFAR-10. We use iterative magnitude pruning with weight rewinding (IMP-WR) and run $10$ iterations of IMP. The unpruned weights are rewound to their values at $2000^{th}$ training step of the original dense network in each iteration. For an initial and quick appreciation, a plot of training loss and test accuracy at different levels of IMP-WR is presented in Fig. 1.
@@
One-shot pruned network is obtained by pruning the weights of the trained dense network $W_{(0)}^{(min\_(0))}$ based on magnitude pruning in one go to attain the desired sparsity [^5] Same sparsity as that of $W_{(10)}^{(min\_(10))}$. and then rewinding the unpruned weights to their values at $W^{(rewind\_point)}$ and retraining. We represent the solution obtained using one-shot pruning by $W^{(one\_shot)}_{(10)}$.
Fine-tuned network is obtained by pruning $20\%$ smallest magnitude weights from $W_{(9)}^{(min\_(9))}$ and then re-training the unpruned weights (without rewinding) with a learning rate of $0.001$ for $40$ epochs. We represent the solution obtained using fine-tuning by $W^{(FT)}_{(10)}$.
Randomly initialized pruned network is obtained by pruning $20\%$ smallest magnitude weights from $W_{(9)}^{(min\_(9))}$ and then randomly initializing the unpruned weights and retraining. We call this network Randomly Initialized Pruned Network and represent the solution by $W^{(RIPN)}_{(10)}$.
Randomly pruned network is obtained by randomly pruning $20\%$ weights from $W_{(9)}^{(min\_(9))}$ and then rewinding and retraining. We call this network Randomly Pruned Network and represent the solution by $W^{(RPN\_1)}_{(10)}$. We also consider another randomly pruned network by randomly pruning the weights of $W_{(0)}^{(min\_(0))}$ in one go to attain the desired sparsity and then rewinding and re-training. We represent the solution obtained in this manner by $W^{(RPN\_2)}_{(10)}$. Note that the only difference between $W^{(RPN\_2)}_{(10)}$ and $W^{(one\_shot)}_{(10)}$ is that while obtaining $W^{(RPN\_2)}_{(10)}$, the weights of $W_{(0)}^{(min\_(0))}$ are pruned randomly whereas while obtaining $W^{(one\_shot)}_{(10)}$, the weights of $W_{(0)}^{(min\_(0))}$ are pruned using magnitude based pruning.
A comparison of training loss and test accuracy between $W_{(10)}^{(min\_(10))}$, $W^{(one\_shot)}_{(10)}$, $W^{(FT)}_{(10)}$, $W^{(RIPN)}_{(10)}$, $W^{(RPN\_1)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$ presented in Fig. 2 shows that $W_{(10)}^{(min\_(10))}$ outperforms the other solutions. To explain the reason behind this, we study the loss landscape of these networks.
Figure 2: Comparison of training loss and test accuracy between $W_{(10)}^{(min\_(10))}$, $W^{(one\_shot)}_{(10)}$, $W^{(FT)}_{(10)}$, $W^{(RIPN)}_{(10)}$, $W^{(RPN\_1)}_{(10)}$ and $W^{(RPN\_2)}_{(10)}$. Left: Training loss. Right: Test accuracy.
@@ IV Results and Findings
IV-A Result 1: Special solutions with small volume exist.
@@
We demonstrate that there also exist another kind of solutions  [^6] We can not say for sure that these solutions are true minima, but they lie in the neighbourhood of minima because the gradient of the vast majority of parameters at these points is zero. which have good generalization performance but have a (relatively) small volume. This implies that volume is not the only criterion for generalization performance; there is more to it. A careful experimental study leads us to hypothesize that the small volume of these solutions is due to very sharp curvature in certain dimensions, but the coefficients in these dimensions are zero. Their volume measure tends to increase when these inferior dimensions are removed (possibly via pruning at another point) but is small when considered in the original space. This makes these solutions almost undiscoverable by SGD in the original space but can be readily discovered in the pruned space. This is the main result of our study.
@@
Despite this at level $(L)$, SGD converges to $W_{(L)}^{(min\_(L))}$ and not to $W^{Pr{(min\_(L-1))}}_{(L)}$. In order to find the underlying reason, we plot the trajectory of SGD for level $L$ and that for level $(L-1)$ projected on level $L$ in Fig. 4.
Figure 4: Comparison of logarithm of training loss versus epoch between level $(L)$ and level $(L-1)$ projected on level $(L)$ for $L$ ranging from $1$ to $10$.
@@
TABLE I: Comparison of inverse volume of basin, $V^{{}^{\prime}}(100)$ at $W^{Pr{(min\_(L-1))}}_{(L)}$ and $W_{(L)}^{(min\_(L))}$.
$L$ Solution $V^{{}^{\prime}}(100)$
1 $W^{Pr{(min\_(0))}}_{(1)}$ 409.242
$W_{(1)}^{(min\_(1))}$ 407.417
2 $W^{Pr{(min\_(1))}}_{(2)}$ 402.678
$W_{(2)}^{(min\_(2))}$ 399.680
3 $W^{Pr{(min\_(2))}}_{(3)}$ 400.855
$W_{(3)}^{(min\_(3))}$ 386.948
4 $W^{Pr{(min\_(3))}}_{(4)}$ 392.829
$W_{(4)}^{(min\_(4))}$ 375.590
5 $W^{Pr{(min\_(4))}}_{(5)}$ 417.984
$W_{(5)}^{(min\_(5))}$ 372.753
6 $W^{Pr{(min\_(5))}}_{(6)}$ 420.431
$W_{(6)}^{(min\_(6))}$ 373.472
7 $W^{Pr{(min\_(6))}}_{(7)}$ 446.758
$W_{(7)}^{(min\_(7))}$ 375.401
8 $W^{Pr{(min\_(7))}}_{(8)}$ 432.505
$W_{(8)}^{(min\_(8))}$ 382.051
9 $W^{Pr{(min\_(8))}}_{(9)}$ 488.004
$W_{(9)}^{(min\_(9))}$ 394.166
10 $W^{Pr{(min\_(9))}}_{(10)}$ 525.621
$W_{(10)}^{(min\_(10))}$ 408.810
We also confirmed this by using the method given by Huang et al. [26] to calculate the radii of the basin. We calculated the radii of the basin along $500$ different random directions around $W_{(10)}^{(min\_(10))}$ and $W^{Pr{(min\_(9))}}_{(10)}$. The radius of the basin in each direction has been calculated by choosing a suitable cut-off value in proximity to the loss values at the minima:
$cut-off=2\times max\{Loss(W_{(10)}^{(min\_(10))}),Loss(W^{Pr{(min\_(9))}}_{(10)})\}$ (5)
This normalized cut-off value turns out to be $1.0$ in this case using equation $5$. A comparison of radii of the basin, around $W_{(10)}^{(min\_(10))}$ and $W^{Pr{(min\_(9))}}_{(10)}$ is given in Fig. 6 and the average radii of basins around $W_{(10)}^{(min\_(10))}$ and $W^{Pr{(min\_(9))}}_{(10)}$ are given in Table II. It is clear from the figure and the table that the average radius of the basin around $W_{(10)}^{(min\_(10))}$ is larger than that around $W^{Pr{(min\_(9))}}_{(10)}$. Since the neural network loss functions lie in high-dimensional spaces, even a small difference in the radii of basins around minima translates to exponentially large disparities in the volume of their surrounding basins.
@@
TABLE III: Comparison of $V^{{}^{\prime}}(100)$ at $W^{RPr{(min\_(L))}}_{(L-1)}$ and $W_{(L-1)}^{(min\_(L-1))}$.
$L$ Solution $V^{{}^{\prime}}(100)$
1 $W^{RPr{(min\_(1))}}_{(0)}$ 418.034
$W_{(0)}^{(min\_(0))}$ 415.42
2 $W^{RPr{(min\_(2))}}_{(1)}$ 415.826
$W_{(1)}^{(min\_(1))}$ 407.417
3 $W^{RPr{(min\_(3))}}_{(2)}$ 403.674
$W_{(2)}^{(min\_(2))}$ 399.68
4 $W^{RPr{(min\_(4))}}_{(3)}$ 394.347
$W_{(3)}^{(min\_(3))}$ 386.948
5 $W^{RPr{(min\_(5))}}_{(4)}$ 391.093
$W_{(4)}^{(min\_(4))}$ 375.59
6 $W^{RPr{(min\_(6))}}_{(5)}$ 391.471
$W_{(5)}^{(min\_(5))}$ 372.753
7 $W^{RPr{(min\_(7))}}_{(6)}$ 394.086
$W_{(6)}^{(min\_(6))}$ 373.472
8 $W^{RPr{(min\_(8))}}_{(7)}$ 398.559
$W_{(7)}^{(min\_(7))}$ 375.401
9 $W^{RPr{(min\_(9))}}_{(8)}$ 411.997
$W_{(8)}^{(min\_(8))}$ 382.051
10 $W^{RPr{(min\_(10))}}_{(9)}$ 427.475
$W_{(9)}^{(min\_(9))}$ 394.166
This is confirmed by a comparison of radii of the basin around $W_{(9)}^{(min\_(9))}$ and $W^{RPr{(min\_(10))}}_{(9)}$ using the method given by Huang et al. [26]. Here, the normalized cut-off value for the radii of the basins turns out to be $0.1$ using equation $5$. A comparison of radii of the basin around $W_{(9)}^{(min\_(9))}$ and $W^{RPr{(min\_(10))}}_{(9)}$ is given in Fig. 8, and the average radii of the two basins are given in Table IV. It is clear from the figure and the table that the average radius of the basin for $W_{(9)}^{(min\_(9))}$ is larger than that for $W^{RPr{(min\_(10))}}_{(9)}$.
@@
To sum up the discussion above, SGD converges to $W_{(L)}^{(min\_(L))}$ at level $(L)$ instead of $W^{Pr{(min\_(L-1))}}_{(L)}$ because the volume of the basin around $W_{(L)}^{(min\_(L))}$ is much larger than the volume of the basin around $W^{Pr{(min\_(L-1))}}_{(L)}$, and the path to $W_{(L)}^{(min\_(L))}$ is steeper than the path to $W^{Pr{(min\_(L-1))}}_{(L)}$. However, in the $(L-1)$ space, the volume of the basin around $W^{RPr{(min\_(L))}}_{(L-1)}$ is much smaller than the volume of basin around $W_{(L-1)}^{(min\_(L-1))}$. This means that the volume comparison gets flipped in the two spaces, $(L-1)$ space and $(L)$ space. To make this more clear, let us consider two spaces, for example, level 00 space and level $1$ space. In level 00 space, the volume of the basin surrounding level 00 solution is larger than the volume of the basin surrounding level $1$ solution; however, in level $1$ space, the volume of the basin surrounding level 00 solution is smaller than the volume of the basin surrounding level $1$ solution. This is shown in Table V.
TABLE V: Comparison of $V^{{}^{\prime}}(100)$ between $W_{(0)}^{(min\_(0))}$ and $W_{(1)}^{(min\_(1))}$ in level 00 and level $1$ space.
Space Solution $V^{{}^{\prime}}(100)$
Level 00 $W_{(0)}^{(min\_(0))}$ (Level 00 solution) 415.42
$W^{RPr{(min\_(1))}}_{(0)}$ (Level $1$ solution in level 00 space) 418.034
Level $1$ $W_{(1)}^{(min\_(1))}$ (Level $1$ solution) 407.417
$W^{Pr{(min\_(0))}}_{(1)}$ (Level 00 solution in level $1$ space) 409.242
IV-B Result 2: Why does the initialization proposed by the lottery ticket hypothesis work well?
The above results also provide a strong clue to the positive role of the initialization proposed by the lottery ticket hypothesis in finding good minima in the sparser weight space. Pruning out the smaller weights leads to a new minimum or a saddle point with a slightly larger training loss value and a smaller volume. Retraining from this point via learning rate rewinding or from a point in close proximity (via weight rewinding) leads SGD to converge to a new minimum, which has a larger volume (compared to the earlier minimum) and a lower training loss. This minimum was not discoverable earlier since it spanned some dimensions where the loss function was steeply increasing and therefore, had an overall smaller volume, and pruning out these dimensions exposed this minimum to the SGD. Hence, the starting minimum acts as a baseline, avoidance of which leads to a better minimum. Choosing any other initialization point places the SGD at a point outside the loss sublevel set and there is no guarantee of finding a better minimum. Fig. 9 presents the training loss along a straight line connecting $W^{(min\_(9))}_{(9)}$ (baseline for $W^{(RIPN)}_{(10)}$) and $W^{(RIPN)}_{(10)}$.
Figure 9: Training loss along a straight line between $W^{(min\_(9))}_{(9)}$ and $W^{(RIPN)}_{(10)}$.
Further, a comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$ (Fig. 10) shows that $W^{(RIPN)}_{(10)}$ has larger eigen values than $W^{(min\_(10))}_{(10)}$, which indicates a smaller volume for the basin around $W^{(RIPN)}_{(10)}$.
Figure 10: Comparison of top-100 positive eigen values of the Hessian at $W^{(RIPN)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$.
IV-C Result 3: Why do we need the Iterative Process, and why does one-shot pruning not work as well?
A comparison of top-100 positive eigen values of the Hessian between $W^{(one\_shot)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$ (Fig. 11) shows that $W^{(one\_shot)}_{(10)}$ has larger eigen values than $W^{(min\_(10))}_{(10)}$, which implies a smaller volume for the basin around $W^{(one\_shot)}_{(10)}$. This brings out and confirms the importance of the iterative process in pruning.
Figure 11: Comparison of top-100 positive eigen values of the Hessian at $W^{(one\_shot)}_{(10)}$ and $W^{(min\_(10))}_{(10)}$.